Suppose an economy's real GDP is $30,000 in year 1 and $31,200 in year 2.what is the growth rate of its real GDP? Asume that population is 100 in year 1 and 102 in year 2. What is the growth rate of GDP per capita?
1
Expert's answer
2021-09-27T10:49:13-0400

Solution:

The growth rate of real GDP:

Economy's real GDP in year 1 = "\$30,000"

Economy's real GDP in year 2 = "\$31,200"


a.). Growth rate of real GDP = "\frac{(31,200 - 30,000)}{30,000} \times 100 = \frac{1,200}{30,000} \times 100 = 4\%"


b.). Population in year 1= 100

So, real GDP per capita in year 1 = "\frac{Real\; GDP}{Population} = \frac{30,000}{100} = \$300"


Population in year 2 = 102

So, real GDP per capita in year 2 = "\frac{Real\; GDP}{Population} = \frac{31,200}{102} = \$305.88"


Growth rate of real GDP per capita = "\frac{(305.88 - 300)}{300} \times 100 = \frac{5.88}{300} \times 100 = 1.96\%"